OpenStudy (anonymous):

Prove: (can only work from the left side) secx+tanx=cosx/1-sinx

OpenStudy (solomonzelman):

You meant "Verify" and work one side independently.

OpenStudy (solomonzelman):

\[Sec~x+Tan~x=\frac{Cos~x}{1}-Sin~x\]

OpenStudy (solomonzelman):

Identities in RED. Proofs in BLACK. \[\color{red} {Secx=1~/~Cosx}\]\[\color{red} {Tan~x=Sin~x~/~Cos~x}\]

OpenStudy (anonymous):

Yes my teacher will only let us work from the left side to prove the right side

OpenStudy (solomonzelman):

\[\frac{1}{Cos~x}+\frac{Sin~x}{Cos~x}=\frac{Cos~x}{1}-Sin~x\]\[\frac{1+Sinx}{Cos~x}=\frac{Cos~x}{1}-Sin~x\]

OpenStudy (solomonzelman):

I am not sure, i have to go

OpenStudy (asnaseer):

@NewK - what have tried so far?

OpenStudy (anonymous):

I got as far as solomonzelman and got stuck there

OpenStudy (asnaseer):

from this step:\[\frac{1+\sin(x)}{\cos(x)}\]try multiplying numerator and denominator by \((1-\sin(x))\)

OpenStudy (anonymous):

Thank you I got it!

OpenStudy (asnaseer):

yw :)